OpenStudy (anonymous):

find the volume of frustum of a regular square pyramid if the base edges are 14 cm and 38 cm, and the measure of one of its lateral edges is 24 cm.

OpenStudy (anonymous):

visualize a frustum of a regular square pyramid. the volume can be expressed as the difference of the volumes of two pyramids: the pyramid with the larger base minus the pyramid with the smaller base that has been cut off.

OpenStudy (anonymous):

Are a of a square-based pyramid is\[\frac{ 1 }{ 3 } side^2 \times height\]

OpenStudy (anonymous):

well i use the apothem formula or simply squaring the base edges, is it the right thing to do?

OpenStudy (anonymous):

I don't know the apothem formula for three-dimensional figures

OpenStudy (anonymous):

i just use the apothem formula of polygons, is it right?

OpenStudy (anonymous):

they tell you it is square-based, just square the side. you have all the information you need to finish the problem.

OpenStudy (anonymous):

yes, if you used the apothem formula it would be right.

OpenStudy (anonymous):

ok so should i consider the lateral edge which is 24 cm to be the slant height of the frustum? so that i can use it to find the altitude by Pythagorean formula

OpenStudy (anonymous):

for the way to solve the problem that I know, you want the height of the imaginary pyramid A and the height of the imaginary pyramid B. There probably is a way to find the volume with the height of the frustum, but I don't know exactly how to do that.

OpenStudy (anonymous):

yes, then how should i use the formula you gave if i don't have the height?

OpenStudy (anonymous):

set up a triangle to represent one side of the imaginary pyramid A.

OpenStudy (anonymous):

ok thats what i did and my final answer is way to far to right one, by the way the formula i am using is\[V = 1/3h(B1+B2+\sqrt{BIB2}\] is this right?

OpenStudy (anonymous):

where B1 is the area of bigger measure and B2 is the area of the smaller one

OpenStudy (anonymous):

where does that formula come from?

OpenStudy (anonymous):

it came from the book

OpenStudy (anonymous):

so what did you use for h?

OpenStudy (anonymous):

the one i get by doing the pythagorean formula which is \[12\sqrt{3}\] from \[h=\sqrt{(24^2)-(19-7)^2}\] is it right?

OpenStudy (anonymous):

http://mathworld.wolfram.com/PyramidalFrustum.html also you can find it here